OpenStudy (anonymous):

Can someone please help me.... simplify. no exponents (m−5)10

OpenStudy (anonymous):

Distributive property of multiplication?

OpenStudy (anonymous):

uh?

OpenStudy (anonymous):

(m-5) * 10 correct?

OpenStudy (anonymous):

that is the same as (m*10) - (5 * 10)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have to add

OpenStudy (anonymous):

Sum/difference of two numbers time a third is the same as the first # times the third add/subtract the second times the third

OpenStudy (anonymous):

the problem is (m ^negative5)10

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[(m ^{-5})^{10}\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

or is it like \[(m ^{-5})10\]

OpenStudy (anonymous):

like the first one

OpenStudy (anonymous):

oh ok well there are a whole set of rules with exponents

OpenStudy (anonymous):

but for this equation, there are rules about exponents being multiplied when it's like that

OpenStudy (anonymous):

yes I know about the product rule, etc but see we haven't had an example like this one thou

OpenStudy (anonymous):

it's like \[m ^{-5} \times m ^{-5} \times m ^{-5} \times m ^{-5} \times m ^{-5} \times m ^{-5} \times m ^{-5} \times m ^{-5} \times m ^{-5} \times m ^{-5}\]

OpenStudy (anonymous):

ohhh pretty much 10 times

OpenStudy (anonymous):

yeah you multiply the exponents

OpenStudy (anonymous):

but what it means when it saids no exponents tho

OpenStudy (anonymous):

so I will multiply 58585....10 times

OpenStudy (anonymous):

opsss 5*5*5*, I meant

OpenStudy (anonymous):

well you can't really get an answer without exponents

OpenStudy (anonymous):

unless you know m

OpenStudy (anonymous):

so you can't simplify this problem

OpenStudy (anonymous):

unless it literally means.. you will just write.. m fifty times

OpenStudy (anonymous):

lol I understand

OpenStudy (anonymous):

thank you sooooo much, this problems are very tricky

OpenStudy (anonymous):

\[\frac{ 1 }{ m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m \times m}\]

OpenStudy (anonymous):

like that..

OpenStudy (anonymous):

which is dumb in my opinion.. because you represent the same thing with exponents

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

can you help me in this one, I did it already but I don't know if I did it correct

OpenStudy (anonymous):

k

OpenStudy (anonymous):

a+7 over7 ;a

OpenStudy (anonymous):

I did this: 9 over 7 + 7 over 7 = 16 over 7

OpenStudy (anonymous):

is it right???

OpenStudy (anonymous):

\[\frac{ (a + 7) }{ 7 } \times a\]

OpenStudy (anonymous):

is that the equation

OpenStudy (anonymous):

well it don't have the x sign, it has ;

OpenStudy (anonymous):

does it mean the same?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok, I added

OpenStudy (anonymous):

Oh how much I hate math :<(

OpenStudy (anonymous):

is it an equal sign

OpenStudy (anonymous):

IDK it only shows the semicolon sign ;

OpenStudy (anonymous):

what are you trying to find

OpenStudy (anonymous):

this problem A+7 over 7 then ;A on the side, the way you wrote it on the side but without the X, only ;

OpenStudy (anonymous):

hmm..

OpenStudy (anonymous):

interesting.... since i'm not sure by what the semicolon means i don't really know what to do